OpenStudy (anonymous):

A stack of nickels and dimes is worth $2.55. find the number of nickels if the number of dimes is 12 less than twice the number of nickels

OpenStudy (anonymous):

i am sorry please mention the value of 1 nickel and 1 dime so i can solve it i live in a country where i dont use this currency

OpenStudy (anonymous):

okay go it! nickel- 5cents dime 10 cents

OpenStudy (anonymous):

easy!

OpenStudy (anonymous):

\[No of dimes+12=2*No. of nickels \right ????\]

OpenStudy (anonymous):

you know that the value of all the coins is 2.55 dolls. there fore 0.10*(no of dimes)+0.05(no of nickels)=2.55 .............(1) no of dimes= 2* no of nickels -12 according to the question .......(2) let us put this in eq-(1) we get 0.10(2x-12)+0.05x=2.55 where x is the number of nickels solve for x we get x=15 let us put thiis value in (2) we totally get 1)no of dimes =18 2)no of nickels =15 there ya go!